H2C2O4 + 2NaOH -> Na2C2O4 + 2H2O


if n(H2C2O4*2H2O) = 6.44*10-6, what is the amount of NaOH in moles?


1
Expert's answer
2020-08-02T15:49:17-0400

H2C2O4 + 2NaOH "\rightarrow"  Na2C2O4 + 2H2O

According to the stoichiometry of the reaction, the number of the moles of the oxalic acid relates to the number of the moles of sodium hydroxide as:

"n(H_2C_2O_4) = \frac{n(NaOH)}{2}" .

The number of the moles of H2C2O4 equals the number of the moles of H2C2O4*2H2O, as there is one molecule of oxalic acid per hydrate unit.

Therefore, if "n(H_2C_2O_4*2H_2O) = 6.44\u00b710^{-6}" mol, the amount of NaOH in moles is "n(NaOH) = 2\u00b7n(H_2C_2O_4*2H_2O) = 1.29\u00b710^{-5}" mol.